# Effects of particle angularity on granular self-organization

### v1

@@ III Results
To conclude, we have studied the effects of grain angularity on self-organisation in quasi-static dynamics of 2D granular assemblies. We simulated regular polygonal grains with gradually sharper corners to test which of the indicators observed in disc systems survive increasing angularity. We have found that some indicators are independent of angularity while others are not. Specifically, as in disc systems, individual cell stresses are well aligned with their orientations. Similarly, the PDFs of the scaled cell stress ratios, $P\left(\hat{h}=h/\bar{h}_{k}\right)$, collapse to a master Weibull form for all cell orders $k$. However, angularity affects the sensitivity of these PDFs to friction; as $N$ decreases, the collapsed $P\left(\hat{h}\right)$ appear to shift systematically toward lower values of $\hat{h}$ as $\mu$ decreases. While one expects cells in low friction assemblies to be less stable, the reason for this difference is unclear. Since the angular corners are expected to have similar traction as in discs, we conjecture that the difference must stem from increased sensitivity to sliding of contacts between flat surfaces. It would be interesting to test this conjecture in future studies.
Another intriguing finding is that the steady-state behaviour changes around $\mu=0.5$: the approach to steady state is slower (Fig. 5 in the SM SM ), there appears a peak in the mean ratio $\sigma_{1}/\sigma_{2}$, and the resistance to shear increases. These phenomena become more noticeable with increasing angularity. These observations suggest an intricate competition between the sliding and rolling mechanisms that govern re-organisation. The probability of surface-to-surface contacts increases with angularity, increasing the sliding probability but reduces the rolling probability. As $\mu$ increases, the probability of rolling increases, but the barrier to rolling is higher for high-angularity grains. Nevertheless, once rolling takes place, the displacements of centers of high-angularity grains are larger. This competition gives rise to optimal stability at an intermediate value of $\mu$,which happens to be around $0.5$ in our bi-disperse polygonal systems.
@@ Acknowledgements.
D. Krengel, T. Matsushima and R. Blumenfeld are grateful for the support of the Grant-in-Aid for Scientific Research 21KK0071 from the Japan Society for the Promotion of Science (JSPS). D. Krengel and T. Matsushima are further grateful for the support of the Grant-in-Aid for Scientific Research 21H01422 from JSPS.
Supplementary Material
V Supplementary figures
V.1 Visualization of the cells
In Fig. 11, the left panel shows examples of cells within the aggregates for different combinations of $N$ and $\mu$. When $\mu=10.0$, angular particles readily form cells with $k>8$. The particles within these cells are rattlers, as the superimposed force network (black lines) illustrates in the right panel.
V.2 The rate of change of $P(k|N)$
Figure 12: The rate of change of the cell conditional probabilities, $dP(k|N)/dt$, with strain for different combinations of friction $\mu$ and angularity $N$.
Fig. 12 illustrates the convergence of the cell conditional probability distributions to their steady-state values, $\mathrm{d}P(k|N)/\mathrm{d}t\to 0$, for different values of $N$ and $\mu$. The fluctuations around the steady state increase somewhat with $\mu$, which is expected to be a general feature. In contrast, the difference between the initial convergence of the very low-friction systems ($\mu=0.01$) and the other systems is attributed to the initial system preparation and the particular procedure we used to generate the steady state.
V.3 $N$-dependence of the collapsed $P(\hat{h})$
The best fits for the order-independent collapse of the PDFs over different $\mu$ for any given $N$, have been shown in the main text to have the Weibull form,
$P\left(\hat{h}\right)=\frac{m}{\lambda}\left(\frac{\hat{h}}{\lambda}\right)^{m-1}e^{-(\bar{h}/\lambda)^{m}}\ .$ (4)
With the exception of hexagonal shapes ($N=6$) the parameters $m$ and $\lambda$ are independent of $N$, as shown in Fig. 13. The PDFs $P\left(\hat{h}\right)$ can also be approximated with a k-Gamma distribution,
although the quality of each fit is notably worse than for the Weibull distribution, and the parameters no longer appear to be independent of $N$, but rather increase as $N$ grows larger.
Figure 13: $N$-dependence of the parameters of the Weibull distribution fit to the cell probability distributions in Fig. 4 in the main text. For reference we also plot the parameters of a k-Gamma distribution to the same data. The error for the Weibull parameters is too small to be seen at this scale.
V.4 The $\mu$-dependence of the steady-state stress ratio
Figure 14: Dependence of the steady state stresses ratio, shown in Fig. 5 in the main file, on $\mu$ and $N$.
The steady-state values of the stress ratio $\sigma_{1}/\sigma_{2}$, shown in Fig. 5 of the main text, show a non-monotonic dependence on the inter-particle friction coefficient, see Fig,14. We observe an increase in $\sigma_{1}/\sigma_{2}$ with $\mu$ until a maximum between $0.1\leq\mu\leq 0.5$. For $\mu\geq 1.0$ the stress ratio decreases again, with the decrease appearing to depend on both cell order and particle shape. As explained in the main text, we understand the reason for this behaviour as a competition between sliding and rolling that leads to an optimal stability at around $\mu=0.5$.
V.5 Global cell orientation distribution
In Fig. 15, we plot the steady-state PDF of cell orientations for all $k>3$. As expected, cells tend to align with the applied load, regardless of $N$. At low friction ($\mu\leq 0.1$), the PDFs approach the uniform form both because cells are less elongated and because the media around cells are more mobile and therefore allow easier cell reorientation.

### v2

@@ IV Conclusions
[…]
Another intriguing finding is that the steady-state behaviour changes around $\mu=0.5$: the approach to steady state is slower (Fig. 9), there appears a peak in the mean ratio $\sigma_{c1}/\sigma_{c2}$, and the resistance to shear increases. These phenomena become more noticeable with increasing angularity. These observations suggest an intricate competition between the sliding and rolling mechanisms that govern re-organisation. The probability of surface-to-surface contacts increases with angularity, increasing the sliding probability but reduces the rolling probability. As $\mu$ increases, the probability of rolling increases, but the barrier to rolling is higher for high-angularity particles. Nevertheless, once rolling takes place, the displacements of centres of high-angularity particles are larger. This competition gives rise to optimal stability at an intermediate value of $\mu$,which happens to be around $0.5$ in our bi-disperse polygonal systems.
In principle, our analysis can be extended to three dimensions (3D). However, to this end some hurdles need to be overcome. While contact data can be readily generated and algorithms to identify 3D cells exist [53], classification of cells into families is significantly more complex. Cells are polyhedra that are characterized by a number of parameters, as opposed to the one order parameter in two dimensions. Each cell $c$ is surrounded by two types of polygonal faces, $N^{c}_{p}$, which are shared with the particles surrounding it, and $N^{c}_{t}$, which constitute throats to nearest-neighbour cells. To characterise cell $c$, we then need to specify: $N^{c}_{p}$; $N^{c}_{t}$; the specific realisation of the order distribution of each of the $N^{c}_{p}$ faces, $P_{p}\left(k\right)$ ($k=1,2,…,N^{c}_{p}$); and the specific realisation of the order distribution of the $N^{c}_{t}$ throat faces, $P_{t}\left(n\right)$ ($n=1,2,\dots,N^{c}_{t}$). A family of cells family not only has the same $N^{c}_{p}$ and $N^{c}_{t}$ faces but it also has the exact same realisation of orders in each type. Given that cells could have 12-14 faces, the number of different combinations of having particle and throat faces is larger than 7500. Even assuming that the orders are distributed between 3 and 9 (the average must be 6), the number of face order realisations for each combination of these is $\mathcal{O}\left(10^{5}\right)$}. Moreover, for good statistics, one would need to have $10^{2}-10^{3}$ specimens of each family member, which means running $10^{7}-10^{8}$ different systems. This is not feasible with current resources.
CRediT authorship contribution statement
DK: Writing – original draft, Writing – review & editing, Visualization, Validation, Software, Methodology, Investigation, Formal analysis, Data curation, Conceptualization. HJ: Writing – review & editing, Software, Visualization, Methodology. RB:: Writing – original draft, Writing – review & editing, Data curation, Conceptualization, Supervision TM: Conceptualization, Supervision, Resources.
@@ Acknowledgements.
DK, TM, and RB are grateful for the support of the Grant-in-Aid for Scientific Research 21KK0071 from the Japan Society for the Promotion of Science (JSPS). DK and TM are also grateful for the support of the Grant-in-Aid for Scientific Research 21H01422 from JSPS.
Appendix A Polygonal DEM method
The polygonal DEM used in this work is based on the monograph by Matuttis and Chen [48]. The interaction force between polygons is based on the particle overlap area $A$ as a stand-in for the physical deformation of two particles (P1 and P2) during contact. We use the centroid $\mathbf{P}$ of the overlap area as force point. The intersection points define the tangential direction $\bm{t}$, which also fixes the normal direction of the contact, $\mathbf{n}$. The elastic normal force
$F_{\mathrm{el}}=\frac{YA}{l},$ (7)
with the Youngs modulus $Y$ and a characteristic length
$l=\frac{4|\mathbf{r}_{\mathrm{A}}||\mathbf{r}_{\mathrm{B}}|}{|\mathbf{r}_{\mathrm{A}}|+|\mathbf{r}_{\mathrm{B}}|}$ (8)
with the contact vectors $\mathbf{r}_{\mathrm{P1,P2}}$ from the the centres of mass to $\mathbf{P}$. We further have a dissipative normal force
$F_{\mathrm{diss}}=\gamma\sqrt{mY}\frac{\dot{A}}{l},$ (9)
with the reduced mass $1/m=1/m_{\mathrm{A}}+1/m_{\mathrm{B}}$ and a damping constant $\gamma$. The total normal force is
$F_{\mathrm{N}}=F_{\mathrm{el}}+F_{\mathrm{diss}}.$ (10)
Friction acting in tangential direction is based on the Cundall-Strack model [54],
$F_{\mathrm{T}}(t)=\left\{\begin{array}[]{ll}F_{\mathrm{T}}(t-\Delta t)-k_{\mathrm{T}}\Delta tv_{\mathrm{T}},&\lvert F_{\mathrm{T}}(t)\rvert\leq\mu F_{\mathrm{N}}\\ \mathrm{sgn}(F_{\mathrm{T}}(t-\Delta t))\mu F_{\mathrm{N}},&\lvert F_{\mathrm{T}}(t)\rvert>\mu F_{\mathrm{N}}\end{array}\right$., (11)
with the relative tangential velocity $v_{\mathrm{T}}$, the tangential stiffness of a spring $k_{\mathrm{T}}$, and $\mu$ as the inter-particle friction coefficient. The total force $\mathbf{F}=F_{\mathrm{N}}\mathbf{n}+F_{\mathrm{T}}\mathbf{t}$ then induces a torque on the particles,
$\tau_{P1,P2}=\mathbf{r}_{\mathrm{P1,P2}}\times\mathbf{F}.$ (12)